OpenStudy (lyssburleson):

Can anyone give me a clear example of Adding/Subtracting and simplifying rational expressions?

OpenStudy (helpmeee):

dropping out

zepdrix (zepdrix):

Rational expressions like this?\[\large\rm \frac{1}{x^2+3x+2}+\frac{x+2}{x+1}\]

OpenStudy (lyssburleson):

yes @zepdrix

zepdrix (zepdrix):

We can at least do this example I guess, and see if you find anything confusing. \(\large\rm x^2+3x+2\) can be factored. Hmm need `factors of 2` which will also `sum to 3`. Oh oh 2*1 = 2 and 2+1 = 3. Ok so we found our values! 2 and 1.\[\large\rm x^2+3x+2=(x+2)(x+1)\]

zepdrix (zepdrix):

So our denominator on the left becomes,\[\large\rm \frac{1}{(x+2)(x+1)}+\frac{x+2}{x+1}\]

zepdrix (zepdrix):

We need a common denominator, notice the denominators already share something in common,\[\large\rm \frac{1}{(x+2)\color{orange}{(x+1)}}+\frac{x+2}{\color{orange}{(x+1)}}\]

zepdrix (zepdrix):

So it appears our second fraction needs an (x+2), both in numerator and denominator,\[\large\rm \frac{1}{(x+2)\color{orange}{(x+1)}}+\frac{x+2}{\color{orange}{(x+1)}}\frac{(x+2)}{(x+2)}\]

zepdrix (zepdrix):

From here, since they have the same denominator, we can rewrite this as a single fraction,\[\large\rm \frac{1+(x+2)(x+2)}{(x+2)(x+1)}\]

zepdrix (zepdrix):

And simplify from there. What do you think? Confusing? :d

OpenStudy (lyssburleson):

Not really, that helped a lot, thank you!